Grade 10Magical Mathematics[Interesting Approach]

In

triangleABC, altitude AD = 18, median BE = 9√5

and

median

CF = 15. Find BC

.

ajinkya bhole
16 Years agoGrade 10

1 Answer

Latika Leekha
12 Years ago
In the given triangle ABC, let AH be the median for BC. Let G be the centroid i.e. G is the point of intersection of the medians AD, BE and CF.
Let GH be the height of \Delta GBC and FK be the height of \DeltaFBC relative to the side BC.
Using the property of median that it divides in the ratio 1:2, we get
CG = 2/3 CF = 2/3. 15 = 10.
BG = 2/3 BE = 2/3 . 9√5 = 6√5.
This means \DeltaBFK is ismilar to \Delta BAD.
Hence, FK/AD = BF/BA.
Now BF= ½ BA (Since CF is the median)
Hence, FK = ½ AD = ½ . 18 = 9.
On similar lines we get, \DeltaCGH is similar to \DeltaCFK.
So, CG/CF = GH/FK
Hence, 10/15 = GH/FK
This means, GH = 2/3 FK = 2/3. 9 = 6.
Now, usinh pythagoras theorem in \DeltaGHB, we get
GH2 + HB2 = GB2.
Hence, 62 + HB2 = (6√5)2.
Hence, this gives, HB2 = 180-36 = 144
Hence, HB = 12.
Again using Pythagoras theorem in \DeltaGHC, we get
GH2 + HC2 = GC2.
62 + HC2 = 102.
Hence, HC = 8.
Now, BC = HB + HC = 12 + 8 = 20.
Hence, the length of BC is 20.